Find the net change in the value of the function between the given inputs.
g(t) = 3 - t^2 ; from -5 to 6

Found 2 solutions by josgarithmetic, math_tutor2020:
Answer by josgarithmetic(39618)
Answer by math_tutor2020(3817)


Plug in t = -5
g(t) = 3 - t^2
g(-5) = 3 - (-5)^2
g(-5) = 3 - 25
g(-5) = -22

Plug in t = 6
g(t) = 3 - t^2
g(6) = 3 - 6^2
g(6) = 3 - 36
g(6) = -33

Subtract the results
g(6) - g(-5) = -33 - (-22) = -33 + 22 = -11

The net change is -11.
It means the output has gone down by 11 units when going from g(-5) = -22 to g(6) = -33.
Use a vertical number line to see why this is the case.

You can think of -22 as being 22 feet below sea level, aka 22 feet underwater.
Move 11 feet further down and you will arrive at the marker -33 which represents being 33 feet underwater.